A person with mass mp = 79 kg stands on a spinning platform disk with a radius of R = 1.83 m and mass md = 183 kg. The disk is initially spinning at ω = 1.8 rad/s. The person then walks 2/3 of the way toward the center of the disk (ending 0.61 m from the center).

5)

What is the centripetal acceleration of the person when she is at R/3?
Answer #1

from conkervation songwon monentur (I + I) W, = ( 2 + I) we (MR4m) Mames) (11 12 + m ) W, = ( 14 2 + y ) wr (M + ) (Man De